OpenStudy (anonymous):

Consider three identical metal spheres, A, B, and C. Sphere A carries a charge of -10q. Sphere B carries a charge of -3q. Sphere C carries no net charge. Spheres A and B are touched together and then separated. Sphere C is then touched to sphere A and separated from it. Lastly, sphere C is touched to sphere B and separated from it. How much charge ends up on sphere C?

OpenStudy (shamim):

total charge of sphere A nd B is -13C. after separation the charge of sphere A will b equal the charge of sphere B. so the charge of sphere A is -6.5C. sphere C has no charge. so after touching A nd C . they will get equal charge. so charge of sphere C is -3.25C

OpenStudy (shamim):

aftet touching with B nd C. sphere C will get total charge =(-3.25-6.5)/2=-9.75/2=-4.875

OpenStudy (shamim):

@jwright16